OpenStudy (anonymous):

How do I factorize ( break down into (x+...) (x-..) or such as those) the following: x^3 + 10

OpenStudy (mathmale):

It might be easier to explain this with a simpler problem. How would you factorize x^3 - 8? I'd change that 8 to 2^3; this is a perfect cube. Then, x^3 - 2^3 = (x - 2)( ????? ) can you fill this in? Now for x^3 + 10: Simply rewrite 10 as a cube:\[10=(\sqrt[3]{10)}^3\]

OpenStudy (mathmale):

Then your x^3 + 10 becomes\[x^3 + (\sqrt[3]{10} )^{3}\] Last hint: Look up "special products and factors"